OpenStudy (anonymous):

The length os the rectangle is 2ft less than twice its width the primeter s 26ft. help asap please!

OpenStudy (anonymous):

length = 2W-2 width = W 2(2W-2)+2W=26 solve that ; W is the width 2W-2 is the length

OpenStudy (anonymous):

Length= L Width= w Given: L=2w-2 Perimeter= 26 26= 2(L+w) 2[(2W-2) + W)= 26 2[(3W-2)=26 3W-2=13 3W=15 Therefore, width= 5 and Length= 2(5)-2=8 RESULT=> Length= 8ft. Width=5ft.